You just won the grand prize in a national writing contest! As your prize, you will receive $500 a month for 50 months. If you can earn 7 percent on your money, what is this prize worth to you today?
Answer #1
Present value of prize money = Monthly cash flow * Present value of annuity of 1
= $             500 * 43.25986
= $ 21,629.93
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.005833)^-50)/0.005833 i = 7%/12 = 0.005833
= 43.2598643 n = 50
Answer #2

Find the PV of an annuity:

PV = C * (1 - PV factor) / r ; PV factor = 1/(1+r)t

Please note that C is the monthly payment and therefore t is the number of months and r is the monthly rate.

Alternatively, use Excel's PV function -> = PV(rate, periods,...). If using Excel make sure payment, rate and periods are consistent (in this case number of months , monthly payment and monthly rate)


source: Black Board
answered by: Mr. Wassi